OpenStudy (anonymous):

for the following equation y=-12x+5 complete the following ordered pair so that it is a solution to the equation ( , -3)

OpenStudy (immanuelv):

plug in -3 to y then you can use this -3 = -12x+5 2= -12x -1/6 =x hope this helps

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

\[-3=-12x+5\] \[-3-5=-12x\] \[-8=-12x\] \[\frac{-8}{-12}=x\] \[x=\frac{2}{3}\]

OpenStudy (anonymous):

i am fairly sure this is right, even though i see two different answers above. step number one should be to subtract 5 from both sides, not add 5

OpenStudy (anonymous):